# Ground state properties of the Heisenberg-compass model on the square lattice

### v1

@@ VII Summary and Discussion
In this work, we examined the zero-temperature quantum ground state phase diagram and the low-temperature classical and quantum phase diagram of the Heisenberg-compass model on the square lattice. Notably, this model admits a Klein duality which facilitates a mapping of the spin-spin interaction parameters from one set to another. This duality analysis partitions the entire parameter space into six distinct regimes, with three of them being Klein-dual counterparts of the remaining three. As a result, the properties of the model in a parameter regime are related to those in its dual regime. For two of those six regimes, the classical zero temperature ground states consist of two symmetry-related discrete configurations. In the remaining four regimes, the classical ground states display an accidental continuous degeneracy characterized by an $O(2)$ manifold. Using classical Monte Carlo simulations and spin-wave analysis, we analyzed the low-temperature classical phase diagram of this model. This reveals six different ordered phases in the six parameter regimes, with four order by disorder (ObD) phases stemming from thermal fluctuations and two energetically ordered phases. By considering quantum fluctuations via a quantum spin-wave analysis at zero temperature, we find that ObD from quantum fluctuations stabilizes the same ordered states as those derived from the preceding classical methodologies. Furthermore, a calculation of the free energy from quantum spin waves at $T>0$ finds that the combined effect of thermal and quantum fluctuations at low temperatures favors the same states as those selected by the quantum fluctuations alone at zero temperature. Additionally, by investigating the zero-temperature quantum ground state phase diagram using numerical exact diagonalization on small finite clusters, we find identical phase diagram as that obtained from the classical analysis and the quantum spin wave analysis.
@@ VII.1 Perspective on applications to materials
It is of interest to briefly discuss the relevance of our work to real materials. Generically, a magnetic material with its magnetic ions on a square lattice will not precisely correspond to the Heisenberg-compass model, even at the nearest-neighbor level, and can possess additional symmetry-allowed interactions. Explicitly, the space group symmetries of the square lattice allow for an additional bond-independent Ising interaction, $S^{z}_{\bm{r}}S^{z}_{\bm{r}^{\prime}}$, on each nearest-neighbor bond Bertinshaw et al. (2019). Incorporating such a term in the original Heisenberg-compass model of Eq. (1), yields
@@
Importantly, a small $\Delta$ ($\cramped{\Delta\ll J}$ and $\cramped{\Delta\ll K}$), either positive or negative, does not lift any of the aforementioned in-plane accidental classical $O(2)$ degeneracies, and therefore does not qualitatively affect the order by disorder physics of the Heisenberg-compass model. This is a generic statement for any symmetry allowed bilinear spin-exchanges, as is found in other order-by-disorder material candidates, such as Er2Ti2O7 Savary et al. (2012). Adapting the arguments of Ref. [Savary et al., 2012], the classical energy of any of the in-plane, accidentally degenerate states is characterized by an order parameter $\bm{m}\equiv(m_{x},m_{y})$ (listed in Table 1) which transforms as $(m_{x},m_{y})\rightarrow(m_{y},-m_{x})$ under the $C_{4}$ symmetry. It is straightforward to show that the only bilinear energy function that can be constructed from $\bm{m}$ is $\propto|\bm{m}|^{2}$ which enjoys an accidental $O(2)$ symmetry. We thus see that these accidental degeneracies, and thus the order-by-disorder, persists even in the presence of generic symmetry allowed bilinear interactions and should thus be relevant in realistic material-relevant extensions of the Heisenberg-compass model.
@@
Although our results are relevant for understanding the ground states properties of a single-layer Ba2IrO4 [$J$approx 65 meV, $K$approx 3.5 meV, and perturbative $\Delta$Katukuri et al. (2014), it does not explain the experimentally observed magnetic ordering in Ba2IrO4 – Néel order along [110] direction Boseggia et al. (2013) (as opposed to the expected ordering in Regime-I of our study). To explain the experimentally observed ordering, Ref. [Katukuri et al., 2014] argues that, along with nearest-neighbour exchanges, one needs to take into account a subset of the inter-layer exchanges which competes with the intra-layer exchanges. In light of these results and the real material context they speak to, it would be interesting to extend the work presented in this study to consider the competition between energetic and ObD state selection in a model that incorporates the symmetry-allowed in-plane $(J,K,\Delta)$ interactions as well as out-of-plane couplings.
@@ VII.2 Avenues for future work on the Heisenberg-compass model
    Much of the literature on ObD has focused on the classical or semi-classical limit ($\cramped{S\rightarrow\infty}$) and does not readily apply in the more realistic quantum limit (e.g. $S=1/2$). A well-understood and unbiased method to study spin-1/2 systems is exact diagonalization Sandvik (2010). As this method is limited to small systems, observables often exhibit large finite-size effects. Therefore, one avenue to better understand ObD and its dynamical implications in spin-1/2 systems may be to characterize the finite-size manifestations of ObD and understand how they might appear in exact diagonalization calculations. The topic of finite-size signatures of ObD has been little explored Lecheminant et al. (1995); Khatua et al. (2021), warranting further investigation. More practically, understanding of such finite-size signatures could be directly relevant to finite size real quantum magnetic systems, such as small magnetic flakes or molecular magnets Park et al. (2016) and trapped ion quantum simulators Kotibhaskar et al. (2023).
@@ Acknowledgements.
We thank Itamar Aharony, Kristian Tyn Kai Chung, R. Ganesh, Felipe Gómez-Lozada, Alex Hickey, Andreas Läuchli, Daniel Lozano-Gómez, and Natalia Perkins for useful discussions. We acknowledge the use of computational resources provided by Digital Research Alliance of Canada. This research was funded by the NSERC of Canada (M.J.P.G, J.G.R) and the Canada Research Chair Program (M.J.P.G, Tier I). G.C.H acknowledges NSERC of Canada funding support through the USRA and CGS-M programs.
@@ Appendix A Luttinger-Tisza method for determination of the classical ground states
In this appendix, we describe the Luttinger-Tisza method Kimchi and Vishwanath (2014); Luttinger and Tisza (1946); Niggemann et al. (2019) to determine the classical ground states of the model of Eq. (1). We start by rewriting the Hamiltonian in the following way,
@@
The modulus square in Eq. (17) comes from the fact that $\bm{S}_{-\bm{q}}=\bm{S}^{*}_{\bm{q}}$ (spin components are real valued in direct space). For a given $\xi$, let $\lambda_{\rm min}$ be the smallest of the minimum values of $\lambda_{1}(\bm{q}),\,\lambda_{2}(\bm{q}),$ and $\lambda_{3}(\bm{q})$ across the first Brillouin zone (BZ). Then Eq. (17) can be written as
@@ Appendix B Spin-wave analysis
In this appendix, we provide the details of the spin wave analysis of the Heisenberg-compass model [Eq. (1)] in the four parameter regimes [Regime-I, II, IV, V] that exhibit ObD, using the Holstein-Primakoff formalism Auerbach (1998). We present the analysis only in Regime-I and IV, and then use the Klein duality to extend the results to the cases of Regime-II and V. We assume for the purpose of this analysis that there is only one sublattice in the magnetic unit cell in the classical ground state. While this is true for the ground state in Regime-IV (ferromagnetic state), the Néel ground state in Regime-I has two magnetic sublattices. However, we can make a transformation to the spins in Regime-I, changing the Néel state to a ferromagnetic state (one sublattice magnetic ordering) so that our one sublattice spin-wave analysis can be applied to obtain the results in Regime-I as well.
@@ B.1 Regime-I
The advantage of performing this transformation is that we now have only one sublattice describing the ground state, rendering the analysis discussed above to find the spin wave spectrum [Eq. (33)] applicable.
@@
Note that $\mathcal{J}^{00}_{0}(\phi)=-2(2J+K)$. Using these Fourier transformed local exchanges, we can compute the linear spin wave spectrum, $\omega_{\bm{q}}(\phi)$, using Eq. (33) as a function of $\phi$. With this spectrum, we next compute the zero-point energy, $\epsilon_{\rm{Q}}(\phi)=(1/2)\sum_{\bm{q}}\omega_{\bm{q}}(\phi)$ as a function of $\phi$. The zero-point energy is found to have minima at $\phi=0,\pi/2,\pi,3\pi/2$ corresponding to the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions. Thus, these are the states picked by quantum ObD at zero temperature in Regime-I.
@@ B.2 Regime-IV
In this regime, the classical ground states are ferromagnetic states pointing along arbitrary directions in the $\bm{\hat{x}}-\bm{\hat{y}}$ plane. To perform the spin-wave analysis, we start with the same reference ground state as in Eq. (36) and the same local frame convention as in Eq. (37). Using the coupling exchange matrix in the global frame in this regime, shown in Eq. (13), we obtain the following Fourier transformed local exchanges
@@
Using these exchanges, we compute the spectrum using Eq. (33), $\omega_{\bm{q}}(\phi)$, as a function of $\phi$. The zero-point energy is found to have minima at $\phi=0,\pi/2,\pi,3\pi/2$, corresponding to the ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions, resulting in quantum ObD at zero temperature.

### v2

@@ VII Summary and Discussion
In this work, we examined the zero-temperature quantum ground state phase diagram and the low-temperature classical and quantum phase diagram of the Heisenberg-compass model on the square lattice. Notably, this model admits a Klein duality which facilitates a mapping of the spin-spin interaction parameters from one set to another. This duality analysis partitions the entire parameter space into six distinct regimes, with three of them being Klein-dual counterparts of the remaining three. As a result, the properties of the model in a parameter regime are related to those in its dual regime. For two of those six regimes, the classical zero temperature ground states consist of two symmetry-related discrete configurations. In the remaining four regimes, the classical ground states display an accidental continuous degeneracy characterized by an $O(2)$ manifold. Using classical Monte Carlo simulations and spin-wave analysis, we analyzed the low-temperature classical phase diagram of this model. These calculations reveal six different ordered phases in the six parameter regimes, with four order by disorder (ObD) phases stemming from thermal fluctuations and two energetically ordered phases. By considering quantum fluctuations via a quantum spin-wave analysis at zero temperature, we find that ObD from quantum fluctuations stabilizes the same ordered states as those derived from the preceding classical methodologies. Furthermore, a calculation of the free energy from quantum spin waves at $T>0$ finds that the combined effect of thermal and quantum fluctuations at low temperatures favors the same states as those selected by the quantum fluctuations alone at zero temperature. Additionally, by investigating the zero-temperature quantum ground state phase diagram using numerical exact diagonalization on small finite clusters, we find identical phase diagram to the one obtained from the classical analysis and the quantum spin wave analysis.
@@ VII.1 Perspective on applications to materials
It is of interest to briefly discuss the relevance of our paper to real materials. Generically, a magnetic material with its magnetic ions on a square lattice will not precisely correspond to the Heisenberg-compass model, even at the nearest-neighbor level, as it may possess additional symmetry-allowed interactions. Explicitly, the space group symmetries of the square lattice allow for an additional bond-independent Ising interaction, $S^{z}_{\bm{r}}S^{z}_{\bm{r}^{\prime}}$, on each nearest-neighbor bond Bertinshaw et al. (2019). Incorporating such a term in the original Heisenberg-compass model of Eq. (1), yields
@@
Importantly, a small $\Delta$ ($\cramped{\Delta\ll J}$ and $\cramped{\Delta\ll K}$), either positive or negative, does not lift any of the aforementioned in-plane accidental classical $O(2)$ degeneracies, and therefore does not qualitatively affect the order by disorder physics of the Heisenberg-compass model. This is a generic statement for any symmetry allowed bilinear spin-exchanges, as is found in other order-by-disorder material candidates, such as Er2Ti2O7 Savary et al. (2012). Adapting the arguments of Ref. [Savary et al., 2012], the classical energy of any of the in-plane, accidentally degenerate states is characterized by an order parameter $\bm{m}\equiv(m_{x},m_{y})$ (listed in Table 1) which transforms as $(m_{x},m_{y})\rightarrow(m_{y},-m_{x})$ under the $C_{4}$ symmetry. It is straightforward to show that the only bilinear energy function that can be constructed from $\bm{m}$ is $\propto|\bm{m}|^{2}$ which enjoys an accidental $O(2)$ symmetry. We thus see that these accidental degeneracies, and thus the order-by-disorder, persists even in the presence of generic symmetry allowed bilinear interactions of arbitrary range and should thus be relevant in realistic material-relevant extensions of the Heisenberg-compass model.
@@
Although our results are relevant for understanding the ground-state properties of a single-layer Ba2IrO4 [$J$approx 65 meV, $K$approx 3.5 meV, and perturbative $\Delta$Katukuri et al. (2014), they do not explain the experimentally observed magnetic ordering in Ba2IrO4 – Néel order along [110] direction Boseggia et al. (2013) (as opposed to the expected ordering in Regime-I of our study). To explain the experimentally observed ordering, Ref. [Katukuri et al., 2014] argues that, along with nearest-neighbour exchanges, one needs to take into account a subset of the inter-layer exchanges which competes with the intra-layer exchanges. However, it has been argued that in Sr2IrO4 Liu and Khaliullin (2019); Kim et al. (2022) the combined effects of strong spin-orbit coupling and Jahn-Teller distortion acting within a layer are responsible for the experimentally observed ordered moment direction. In light of these results and the similar material context, it would be interesting to extend the work presented in this study to consider the competition between energetic and ObD state selection in a model of Ba2IrO4 incorporating both the symmetry-allowed in-plane $(J,K,\Delta)$ interactions as well as spin-lattice and other couplings discussed in more detail for Sr2IrO4 Liu and Khaliullin (2019); Kim et al. (2022).
@@ VII.2 Avenues for future work on the Heisenberg-compass model
    Much of the literature on ObD has focused on the classical or semi-classical limit ($\cramped{S\rightarrow\infty}$) and does not readily apply in the more realistic quantum limit (e.g. $S=1/2$). A well-understood and unbiased method to study spin-1/2 systems is exact diagonalization Sandvik (2010). As this method is limited to small systems, observables often exhibit large finite-size effects. Therefore, one avenue to better understand ObD and its dynamical implications in spin-1/2 systems may be to characterize the finite-size manifestations of ObD and understand how they might appear in exact diagonalization calculations. The topic of finite-size signatures of ObD has been little explored Lecheminant et al. (1995); Khatua et al. (2021), warranting further investigation. More practically, achieving an understanding of such finite-size signatures could be directly relevant to finite size real quantum magnetic systems, such as small magnetic flakes or molecular magnets Park et al. (2016) and trapped ion quantum simulators Kotibhaskar et al. (2024).
@@ Acknowledgements.
We thank Itamar Aharony, Kristian Tyn Kai Chung, R. Ganesh, Felipe Gómez-Lozada, Alex Hickey, Andreas Läuchli, Daniel Lozano-Gómez and Natalia Perkins for useful discussions. We also thank Giniyat Khaliullin for useful comments on an early version of the manuscript. We acknowledge the use of computational resources provided by Digital Research Alliance of Canada. This research was funded by the NSERC of Canada (M.J.P.G, J.G.R) and the Canada Research Chair Program (M.J.P.G, Tier I). G.C.H acknowledges funding from the NSERC of Canada through the USRA and CGS-M programs.
@@ Appendix A Luttinger-Tisza method for determination of the classical ground states
In this appendix, we describe the Luttinger-Tisza method Kimchi and Vishwanath (2014); Luttinger and Tisza (1946); Niggemann et al. (2019) used to determine the classical ground states of the model of Eq. (1). We start by rewriting the Hamiltonian in the following way,
@@
The modulus square in Eq. (17) comes from the fact that $\bm{S}_{-\bm{q}}=\bm{S}^{*}_{\bm{q}}$ (spin components are real valued in direct space). For a given $\xi$, let $\lambda_{\rm min}$ be the smallest of the minimum values of $\lambda_{1}(\bm{q}),\,\lambda_{2}(\bm{q}),$ and $\lambda_{3}(\bm{q})$ across the first Brillouin zone (BZ). Then, Eq. (17) can be written as
@@ Appendix B Spin-wave analysis
In this appendix, we provide the details of the spin wave analysis of the Heisenberg-compass model [Eq. (1)] in the four parameter regimes [Regime-I, II, IV, V] that exhibit ObD, using the Holstein-Primakoff formalism Auerbach (1998). We present the analysis only in Regime-I and IV, and then use the Klein duality to extend the results to the cases of Regime-II and V. We assume for the purpose of this analysis that there is only one sublattice in the magnetic unit cell in the classical ground state. While this is true for the ground state in Regime-IV (ferromagnetic state), the Néel ground state in Regime-I has two magnetic sublattices. However, we can make a transformation to the spins in Regime-I, changing the Néel state to a ferromagnetic state (one sublattice magnetic ordering) so that one sublattice spin-wave analysis can be applied to obtain the results in Regime-I as well.
@@ B.1 Regime-I
The advantage of performing this transformation is that we now have only one sublattice describing the ground state, making the one-sublattice spin-wave analysis discussed above directly applicable in Regime-I to find the spin-wave spectrum [Eq. (33)].
@@
Note that $\mathcal{J}^{00}_{0}(\phi)=-2(2J+K)$. Using these Fourier transformed local exchanges, we can compute the linear spin wave spectrum, $\omega_{\bm{q}}(\phi)$, using Eq. (33) as a function of $\phi$. With this spectrum in hand, we next compute the zero-point energy, $\epsilon_{\rm{Q}}(\phi)=(1/2)\sum_{\bm{q}}\omega_{\bm{q}}(\phi)$ as a function of $\phi$. The zero-point energy is found to have minima at $\phi=0,\pi/2,\pi,3\pi/2$, which correspond to the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions and these are thus the states picked by quantum ObD at zero temperature in Regime-I.
@@ B.2 Regime-IV
In this regime, the classical ground states are ferromagnetic states pointing along arbitrary directions in the $\bm{\hat{x}}-\bm{\hat{y}}$ plane. To perform the spin-wave analysis, we start with the same reference ground state as in Eq. (36) and the same local frame convention as in Eq. (37). Using the coupling exchange matrix in the global frame in this regime, given by Eq. (13), we obtain the following Fourier transformed local exchanges
@@
Using these exchanges, we compute the spectrum using Eq. (33), $\omega_{\bm{q}}(\phi)$, as a function of $\phi$. The zero-point energy is found to have minima at $\phi=0,\pi/2,\pi,3\pi/2$, corresponding to the ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions, which result from quantum ObD at zero temperature.